A car moves towards north at a speed of 54 km per hour for 1 hr then it moves eastward with same speed for same duration tbe average speed and average velocity of car for complete journey is
 

Grade:12th pass

1 Answers

Agrata Singh
5 years ago
Let north dirxn be denoted as Y axis and east as X axis
Distance travelled north=54x1=54 km
Distance travelled east=54x1=54 km
Total distance travelled =54+54=108km
Displacement s= 54j +54i
                     |s|=54√2km
Average speed=(Total distance)/(total time)= 108/2= 54km/hr
Average Velocity =(total displacement)/(total time)= 54√2/2= 27√2 km/hr